# Magnetic Field Line Chaos, Cantori, and Turnstiles in Toroidal Plasmas

### v2

@@ 5.1 Efficient way to obtain information about cantori and turnstiles
Conventionally subtleties of magnetic field line evolution are understood using Poincaré plots. (Boozer:2025:Efficient) has shown that far more information can be obtained by Fourier decomposing $R(\varphi)$ and $Z(\varphi)$ of magnetic field lines that are followed in ordinary $(R,\varphi,Z)$ cylindrical coordinates to find surfaces on which magnetic field lines lie or near which they linger. Knowing these surfaces, $\vec{B}\cdot\hat{n}$ can be determined, cantori identified, the magnitude and collumation of the magnetic flux that leaks through these surfaces in turnstiles can be found.
@@ 5.2 Subtleties of cantori and turnstiles in non-resonanant divertors
An intuitive understanding of subtleties of cantori and turnstiles can be obtained from a study of the localization of the plasma flows in a stellarator non-resonant divertor. This can be done by following field lines from the point where they enter the plasma chamber until they leave. The strong helical fields outside of the magnetic surfaces of a stellarator cause almost all magnetic field lines to exit the plasma chamber within approximately a stellarator period of where they entered. However, there are narrow helical stripes on the walls, which take up only a few percent of the wall area, with entry points that give lines that make many toroidal transits before restriking the wall. Indeed, there are entry points within these stripes from which field lines make an arbitrarily large number of toroidal transits (Punjabi:2025). The field lines that come closest to the confined plasma take the largest number of toroidal transits between wall entry and departure. These lines form a tube of magnetic flux that can be filled with plasma by cross-field diffusion and produce the divertor. This flux tube is roughly circular but strikes the wall in a helical stripe since the magnetic field component tangential to the stellarator walls is strong in comparison to the normal field.
@@
While the field line is making toroidal transits between the entry and the exit turnstile of the outermost important cantorus, it can intercept an even smaller turnstile of a cantoris that is closer to the last confining magnetic surface. These lines become trapped for an even larger number of toroidal transits before they intercept an existing turnstile. The strong correlation between the closeness with which a field line approaches the outermost confining magnetic surface and the number of toroidal transits the line makes is illustrated in (Garcia:2025) and (Punjabi:2025).
The location of the outermost important turnstile is determined by a strong curvature in the outermost magnetic surface, which are called helical edges. They resemble the X-lines of a tokamak divertor. Unlike X-lines, which are toroidally continuous in axisymmetry and require the rotational transform to be zero, helical edges extend only over a fraction of stellarator period. A sharp helical edge on a magnetic surface, or an X-line of a tokamak divertor, can only be produced by currents for more distant than the radius of curvature of the magnetic surface when the edge is parallel to a magnetic field line. Since the helical edge is not a closed curve, it does not define the rotational transform of the surface. (Bader:2018) has shown the locations on the wall that are struck by the field lines of a non-resonant divertor are robust against plasma changes and are determined by strong variation in the curvature of the outermost magnetic surface.
6 Magnetic reconnection
@@
6.1 Magnetic field line velocity $\vec{u}_{\bot}$
@@
The loop voltage in a region of spatially bounded field lines is defined by Equation (3). When $\vec{\nabla}V_{\ell}=0$ the magnetic evolution is ideal. Faraday’s law then implies the evolution of Clebsch coordinates $\alpha$ and $\beta$ where $\vec{B}=\vec{\nabla}\alpha\times\vec{\nabla}\beta$ is consistent with $d\alpha/dt=\partial\alpha/\partial t+\vec{u}_{\bot}\cdot\vec{\nabla}\alpha=0$ and $d\beta/dt=\partial\beta/\partial t+\vec{u}_{\bot}\cdot\vec{\nabla}\beta=0$. Since $\alpha$ and $\beta$ define a magnetic field line, $\vec{u}_{\bot}$ is the velocity of the magnetic field lines when the gradient of the loop voltage is zero. This is in essence the proof given in (Newcomb:1958). The physical interpretation of $\vec{u}_{\bot}$ is subtle when $\vec{\nabla}V_{\ell}\neq 0$, but $\vec{u}_{\bot}$ is still defined and for simplicity of language can be called the magnetic field line velocity. The fact that $\vec{u}_{\bot}$ depends on the frame of reference that is used, which can change if canonical transformations are made to $(\psi_{t},\theta,\varphi)$ as the system evolves, has created confusion about the existence of $\vec{u}_{\bot}$.
Equation (17) for $\vec{E}$ follows from mathematics alone—the representation of vectors in a three dimensional space. Although (Boozer:2025:Magnetic) has shown that Equation (17) is far more generally valid, the proof is more obvious when $\vec{B}\cdot\vec{\nabla}\varphi\neq 0$. An arbitrary vector $\vec{E}$ in three-space can be represented in terms of another arbitrary vector $\vec{B}$ when all three components of $\vec{E}$ can be represented. The component $\vec{B}\cdot\vec{E}=-\vec{B}\cdot\vec{\nabla}\Phi+V_{\ell}\vec{B}\cdot\vec{\nabla}\varphi$ is equivalent to
@@
(Aref:1984) illustrated the exponentially large effect on mixing in fluids when the velocity with which they are stirred is chaotic. (Tang:1999) used Lagrangian coordinates to show the effective spatial diffusion increases exponentially in time.
6.2 Magnetic reconnection when $\vec{B}$ is nearly ideal
When $\vec{B}$ is ideal and $\vec{u}_{\bot}$ is chaotic, the surfaces that enclose a volume of magnetic flux increase their surface area exponentially without limit even though the enclosed magnetic flux is a constant, Figure 3. (Boozer and Elder 2021) gave an example of smooth large-scale non-turbulent magnetic field line flow $\vec{u}_{\bot}$ that is chaotic. Their example is particularly interesting because it puts no helicity into the system. As they discuss, unlike magnetic energy, the dissipation of magnetic helicity cannot be significantly enhanced by turbulence. In a system like the solar corona in which magnetic flux tubes are driven by plasma motion in the photosphere, helicity buildup cannot be bounded other than by the ejection of the loop from the sun, a coronal mass ejection.
@@ 6.2 Magnetic reconnection when $\vec{B}$ is nearly ideal
6.3 Limitation on reconnection due to electron inertia
The finite electron mass can cause magnetic reconnection even in the absence of any other non-ideal effect. However, (Boozer:2025:Electron) proved a remarkable result. When electron inertia is the only non-ideal effect in the evolution of a magnetic field $\vec{B}$, there is a related field that evolves ideally. This field is $\vec{\mathcal{B}}\equiv\vec{B}+\vec{\nabla}\times\left((c/\omega_{pe})^{2}\mu_{0}\vec{j}\right)$ with $\omega_{pe}$ the plasma frequency and $\vec{j}$ the current density. In three dimensional space, the practical importance of the ideal evolution of $\vec{\mathcal{B}}$ on its reconnection appears limited. Section 6.2 shows that when the evolution velocity of modified field is chaotic, $\vec{\mathcal{B}}$ will reconnect on a timescale that depends only logarithmically on any non-ideal effect that is diffusive, such as resistivity.
6.4 Magnetic surface breakup and electrostatic instabilities
(Nevins:2011), (Connor:2013), and (Terry:2015) have discussed magnetic surface breakup that is produced by what are called electrostatic instabilities, primarily the ion temperature gradient (ITG) instabilities. The focus has been on the enhancement of the electron heat transport, which is not large when the plasma $\beta\equiv 2\mu_{0}p/B^{2}$ is small.
An effect that can arise before the electron heat transport is appreciable is a modification of the radial electric field required to preserve quasi-neutrality in stellarators. This modification is of great practical importance because neoclassical transport in stellarators usually gives more rapid ion than electron transport, which requires an electric field that confines ions—particularly high charge state impurities. (Helander:2008) discussed the lack of effect of electrostatic microturbulence on the large scale radial electric field though the turbulent Reynolds stress can produce zonal flows that tend reduce the microturbulent transport. They did not discuss the effect of the breaking of the magnetic surfaces by the turbulence. The enhanced radial transport of electrons due to broken surfaces does modify the radial electric field, but the required momentum transport to break the surfaces is subtle. As the author has discussed but not yet published, the radial transport of the electrons gives an effective perpendicular viscosity force. (Alcison:2023) found a large expulsion of impurities in W7-X experiments when the microturbulent transport exceeded the expected neoclassical, which requires an explanation.
@@ 6.4 Magnetic surface breakup and electrostatic instabilities
7 Discussion
Once the concepts of magnetic field line chaos, cantori, and turnstiles are understood their importance to determining the behavior of toroidal fusion plasmas becomes clear. This section gives a few additional examples of their applications, but also expresses concern about the future of innovation within U. S. fusion program.
@@ 7.1 Tokamak disruptions
As noted in the Introduction, the current density required to create an exponentiation $\sigma$ is only linearly dependent of $\sigma$, see Appendix A.3, as is the timescale for large scale reconnection, see Section 6. Section 4 showed an ideal perturbation can lead to chaotic field lines through a large volume of the plasma even though perfect magnetic surfaces are preserved. Section 6 shows that time required before reconnection occurs is essentially determined by timescale on which the exponentiation of the chaos grows. The time required for a large scale reconnection has only a logarithmic dependence on resistivity, which makes this time essentially independent of resistivity when the resistivity is small. Shortly after the paper (Boozer:2022) was was published showing this, (Jardin:2022) carried out simulations of disruptions that were extremely fast due to an ideal mode becoming unstable. Their simulations confirmed that the time to disruption was essentially independent of the plasma resistivity to the extent the code could maintain adequate resolution. That is, any large scale mode, resistive or ideal, can produce a disruption on a timescale determined by the growth rate of the mode unless it self-stabilizes at a sufficiently small amplitude. Non-self-stabilizing modes could cause a stellarator to disrupt, but such modes are easier to avoid in stellarators than in tokamaks.
@@
When the magnetic evolution is near ideal, as in a high-temperature tokamak, (Boozer:2025:Magnetic) showed only a small fraction of the energy released by the reconnection is directly dissipated—most goes into Alfvén waves. This is unlike reconnection in two-coordinate systems in which the energy released by the reconnection is damped on the same timescale as the reconnection. (Boozer:2020) showed that shear Alvén waves moving along the reconnecting field lines would cause the flattening of the $j_{||}/B$ profile, which leads to the current spike observed in disruptions. The simulations of (Nardon:2023) were consistent with the predicted timescale, and they saw magnetic fluctuations that appeared to be shear Alfvén waves.
@@
(Heyvaerts:1983) and (Similon:1989) showed that when shear Alfvén waves propagate along chaotic field lines they are quickly damped. (Huang-Bhattacharjee:2022) defined reconnection, even in three-coordinate problem, as the transfer of energy from the reconnecting field to the plasma, not the breaking of field line connections. Their definition, which is common in two-coordinate models, led them to the odd conclusion that the chaos of the field lines was irrelevant to reconnection—only the formation of localized currents, which are caused by Alfvén wave damping, are relevant.
@@
The rapid spreading of impurities across the entire plasma is a commonly observed feature of tokamak disruptions. The impurities greatly reduce the damage produced by disruptions by radiating the plasma thermal energy relatively uniformly over the walls rather than having it concentrated in spots where turnstiles formed within the plasma strike the walls as is often the case with runaway electrons. The reason for the impurity spreading did not have a firm theoretical basis, but the Bohm-like transport that is produced when there is an electron pressure gradient along chaotic field lines, (Boozer:2024:Electric) provides an explanation. The simulations of (Nardon:2023) saw electric fields consistent in magnitude, but the equation in their code which gave the electric field was not simply related to the two-fluid equations that (Boozer:2024:Electric) used to obtain the Bohm-like transport.
@@
(Breizman:2019) reviewed the problem of the conversion of the plasma current from being carried by near-thermal to being carried by relativistic electrons, which is often viewed as the most dangerous feature of disruptions. Even though the total energy in the relativistic electrons is only about ten percent of the pre-disruption plasma thermal energy, impurity radiation spreads the thermal energy losses almost uniformly over the wall. Relativistic electron losses onto the walls can be extremely concentrated in both space and time, which means they are far more destructive. As pointed out by (Boozer-Punjabi:2016) the spatial and temporal concentration can be explained by the breaking of the outermost magnetic surface that is confining the runaway electrons in internal bounded chaotic region. As the outermost surface breaks it forms a cantorus. The slower the turnstiles form compared to the timescale for the relativistic electrons to cover the chaotic region the more tightly collimated are the turnstiles on the cantorus. When the turnstiles form rapidly compared to the timescale for the relativistic electrons to cover the chaotic region, the turnstiles become arbitrary large, which spreads the relativistic electrons so broadly on the walls that they do no damage. The highly damaging case is often observed and brings fear, but (Reux:2021) and (Paz-Soldan:2021) have observed on the JET and the DIII-D tokamaks the benign spreading of the relativistic electrons when the plasma becomes unstable to a rapidly growing magnetic perturbation.
@@
(Boozer:2025:Constraints) has pointed out that the frequency with which disruptions occur in tokamaks may be in large part be due to the far larger poloidal flux produced by the plasma current than the change in the poloidal flux that occurs if the profile of the density of the net plasma current $j_{||}$ is varied over the full range of profiles stable to tearing modes. This flux ratio is approximately a factor of six. Equation (2), $\partial\psi_{p}/\partial t=V_{\ell}$, implies the plasma will disrupt within a shorter time than the time required to shutdown by removing the plasma-produced poloidal flux or the natural length of a flattop pulse unless a spatially constant loop voltage remains consistent with tearing mode stability. As shown in Equation 19, the magnetic field can be held fixed when $\vec{\nabla}V_{\ell}=0$. An implication is that poloidal flux can be removed or maintained without producing a disruption when
@@
The profile of $j_{||}$ can in principle be controlled by plasma heating and current drive. During the flattop of a fusion pulse, the required power to produce a large change in the $j_{||}$ profile is comparable to the power produced by the alpha-particle heating. This is obvious for direct heating. (Boozer:1988) showed the power required to drive the full plasma current in a power plant is greater than $j_{cd}E_{ch}$ integrated over the plasma volume where $E_{ch}$ is the Connor-Hastie electric field (Connor:1975). The implication is the power required to drive the total plasma current is comparable to the power produced by alpha-particle heating. Careful control of the injected power is required to properly control the profile of $j_{||}$, but the available diagnostics may not provide adequate information for such control even if it were in principle possible. It is unclear whether the ratio of the total plasma flux produced by the plasma current versus the change in the poloidal flux if the current profile evolved from disruption stable to unstable has been calculated for the ARC designs of Commonwealth Fusion Systems or what the strategy is for ensuring a stable profile evolution. This disruption issue can be circumvented by the use of stellarator magnetic field to produce the rotational transform by external coil currents rather than by the plasma current.

### v3

@@ 5.1 Efficient way to obtain information about cantori and turnstiles
Conventionally subtleties of magnetic field line evolution are understood using Poincaré plots. (Boozer, 2025b) has shown that far more information can be obtained by Fourier decomposing $R(\varphi)$ and $Z(\varphi)$ of magnetic field lines that are followed in ordinary $(R,\varphi,Z)$ cylindrical coordinates to find surfaces on which magnetic field lines lie or near which they linger. Knowing these surfaces, $\vec{B}\cdot\hat{n}$ can be determined, cantori identified, the magnitude and collumation of the magnetic flux that leaks through these surfaces in turnstiles can be found.
@@ 5.2 Subtleties of cantori and turnstiles in non-resonanant divertors
An intuitive understanding of subtleties of cantori and turnstiles can be obtained from a study of the localization of the plasma flows in a stellarator non-resonant divertor. This can be done by following field lines from the point where they enter the plasma chamber until they leave. The strong helical fields outside of the magnetic surfaces of a stellarator cause almost all magnetic field lines to exit the plasma chamber within approximately a stellarator period of where they entered. However, there are narrow helical stripes on the walls, which take up only a few percent of the wall area, with entry points that give lines that make many toroidal transits before restriking the wall. Indeed, there are entry points within these stripes from which field lines make an arbitrarily large number of toroidal transits (Punjabi and Boozer, 2025). The field lines that come closest to the confined plasma take the largest number of toroidal transits between wall entry and departure. These lines form a tube of magnetic flux that can be filled with plasma by cross-field diffusion and produce the divertor. This flux tube is roughly circular but strikes the wall in a helical stripe since the magnetic field component tangential to the stellarator walls is strong in comparison to the normal field.
@@
While the field line is making toroidal transits between the entry and the exit turnstile of the outermost important cantorus, it can intercept an even smaller turnstile of a cantoris that is closer to the last confining magnetic surface. These lines become trapped for an even larger number of toroidal transits before they intercept an existing turnstile. The strong correlation between the closeness with which a field line approaches the outermost confining magnetic surface and the number of toroidal transits the line makes is illustrated in (Garcia et al., 2025) and (Punjabi and Boozer, 2025).
The location of the outermost important turnstile is determined by a strong curvature in the outermost magnetic surface, which are called helical edges. They resemble the X-lines of a tokamak divertor. Unlike X-lines, which are toroidally continuous in axisymmetry and require the rotational transform to be zero, helical edges extend only over a fraction of stellarator period. A sharp helical edge on a magnetic surface, or an X-line of a tokamak divertor, can only be produced by currents for more distant than the radius of curvature of the magnetic surface when the edge is parallel to a magnetic field line. Since the helical edge is not a closed curve, it does not define the rotational transform of the surface. (Bader et al., 2018) has shown the locations on the wall that are struck by the field lines of a non-resonant divertor are robust against plasma changes and are determined by strong variation in the curvature of the outermost magnetic surface.
6 Applications
Once the concepts of magnetic field line chaos, cantori, and turnstiles are understood their importance to determining the behavior of plasmas becomes clear. This section discusses three araes of application: magnetic reconnection, tokamak disruptios, and stellarator physics.
6.1 Magnetic reconnection
@@
6.1.1 Magnetic field line velocity $\vec{u}_{\bot}$
@@
The loop voltage in a region of spatially bounded field lines is defined by Equation (3). When $\vec{\nabla}V_{\ell}=0$ the magnetic evolution is ideal. Faraday’s law then implies the evolution of Clebsch coordinates $\alpha$ and $\beta$ where $\vec{B}=\vec{\nabla}\alpha\times\vec{\nabla}\beta$ is consistent with $d\alpha/dt=\partial\alpha/\partial t+\vec{u}_{\bot}\cdot\vec{\nabla}\alpha=0$ and $d\beta/dt=\partial\beta/\partial t+\vec{u}_{\bot}\cdot\vec{\nabla}\beta=0$. Since $\alpha$ and $\beta$ define a magnetic field line, $\vec{u}_{\bot}$ is the velocity of the magnetic field lines when the gradient of the loop voltage is zero. This is in essence the proof given in (Newcomb, 1958). The physical interpretation of $\vec{u}_{\bot}$ is subtle when $\vec{\nabla}V_{\ell}\neq 0$, but $\vec{u}_{\bot}$ is still defined and for simplicity of language can be called the magnetic field line velocity. The fact that $\vec{u}_{\bot}$ depends on the frame of reference that is used, which can change if canonical transformations are made to $(\psi_{t},\theta,\varphi)$ as the system evolves, has created confusion about the existence of $\vec{u}_{\bot}$.
Equation (17) for $\vec{E}$ follows from mathematics alone—the representation of vectors in a three dimensional space. Although (Boozer, 2025c) has shown that Equation (17) is far more generally valid, the proof is more obvious when $\vec{B}\cdot\vec{\nabla}\varphi\neq 0$. An arbitrary vector $\vec{E}$ in three-space can be represented in terms of another arbitrary vector $\vec{B}$ when all three components of $\vec{E}$ can be represented. The component $\vec{B}\cdot\vec{E}=-\vec{B}\cdot\vec{\nabla}\Phi+V_{\ell}\vec{B}\cdot\vec{\nabla}\varphi$ is equivalent to
@@
(Aref, 1984) illustrated the exponentially large effect on mixing in fluids when the velocity with which they are stirred is chaotic. (Tang and Boozer, 1999) used Lagrangian coordinates to show the effective spatial diffusion increases exponentially in time.
6.1.2 Magnetic reconnection when $\vec{B}$ is nearly ideal
When $\vec{B}$ is ideal and $\vec{u}_{\bot}$ is chaotic, the surfaces that enclose a volume of magnetic flux increase their surface area exponentially without limit even though the enclosed magnetic flux is a constant. The exponential increase in the flux tube perimeters in Figure 3 correspond to an exponential increase in surface area if a three rather than a two dimensional plot were made. (Boozer and Elder 2021) gave an example of smooth large-scale non-turbulent magnetic field line flow $\vec{u}_{\bot}$ that is chaotic. Their example is particularly interesting because it puts no helicity into the system. As they discuss, unlike magnetic energy, the dissipation of magnetic helicity cannot be significantly enhanced by turbulence. In a system like the solar corona in which magnetic flux tubes are driven by plasma motion in the photosphere, helicity buildup cannot be bounded other than by the ejection of the loop from the sun, a coronal mass ejection.
@@ 6.1.2 Magnetic reconnection when $\vec{B}$ is nearly ideal
6.1.3 Limitation on reconnection due to electron inertia
The finite electron mass can cause magnetic reconnection even in the absence of any other non-ideal effect. However, (Boozer, 2026b) proved a remarkable result. When electron inertia is the only non-ideal effect in the evolution of a magnetic field $\vec{B}$, there is a related field that evolves ideally. This field is $\vec{\mathcal{B}}\equiv\vec{B}+\vec{\nabla}\times\left((c/\omega_{pe})^{2}\mu_{0}\vec{j}\right)$ with $\omega_{pe}$ the plasma frequency and $\vec{j}$ the current density. In three dimensional space, the practical importance of the ideal evolution of $\vec{\mathcal{B}}$ on its reconnection appears limited. Section 6.1.2 shows that when the evolution velocity of modified field is chaotic, $\vec{\mathcal{B}}$ will reconnect on a timescale that depends only logarithmically on any non-ideal effect that is diffusive, such as resistivity.
6.1.4 Magnetic surface breakup and electrostatic instabilities
(Nevins et al., 2011), (Connor, Hastie, and Zocco, 2013), and (Terry et al., 2015) have discussed magnetic surface breakup that is produced by what are called electrostatic instabilities, primarily the ion temperature gradient (ITG) instabilities. The focus has been on the enhancement of the electron heat transport, which is not large when the plasma $\beta\equiv 2\mu_{0}p/B^{2}$ is small.
An effect that can arise before the electron heat transport is appreciable is a modification of the radial electric field required to preserve quasi-neutrality in stellarators. This modification could be of great practical importance because neoclassical transport in stellarators usually gives more rapid ion than electron transport, which requires an electric field that confines ions—particularly high charge state impurities. (Helander and Simakov, 2008) discussed the lack of effect of electrostatic microturbulence on the large scale radial electric field though the turbulent Reynolds stress can produce zonal flows that tend reduce the microturbulent transport. They did not discuss the effect of the breaking of the magnetic surfaces by the turbulence. The enhanced radial transport of electrons due to broken surfaces does modify the radial electric field, but the required momentum transport to break the surfaces is subtle. (Boozer, 2026a) has shown the radial transport of the electrons gives an effective perpendicular viscosity force. (Alcisón et al., 2023) found a large expulsion of impurities in W7-X experiments when the microturbulent transport exceeded the expected neoclassical, which requires an explanation.
@@ 6.2 Tokamak disruptions
As noted in the Introduction, the current density required to create an exponentiation $\sigma$ is only linearly dependent of $\sigma$, see Appendix A.3, as is the timescale for large scale reconnection, see Section 6.1. Section 4 showed an ideal perturbation can lead to chaotic field lines through a large volume of the plasma even though perfect magnetic surfaces are preserved. Section 6.1 shows that time required before reconnection occurs is essentially determined by timescale on which the exponentiation of the chaos grows. The time required for a large scale reconnection has only a logarithmic dependence on resistivity, which makes this time essentially independent of resistivity when the resistivity is small. Shortly after the paper (Boozer, 2022) was was published showing this, (Jardin et al., 2022) carried out simulations of disruptions that were extremely fast due to an ideal mode becoming unstable. Their simulations confirmed that the time to disruption was essentially independent of the plasma resistivity to the extent the code could maintain adequate resolution. That is, any large scale mode, resistive or ideal, can produce a disruption on a timescale determined by the growth rate of the mode unless it self-stabilizes at a sufficiently small amplitude. Non-self-stabilizing modes could cause a stellarator to disrupt, but such modes are easier to avoid in stellarators than in tokamaks.
@@
When the magnetic evolution is near ideal, as in a high-temperature tokamak, (Boozer, 2025c) showed only a small fraction of the energy released by the reconnection is directly dissipated—most goes into Alfvén waves. This is unlike reconnection in two-coordinate systems in which the energy released by the reconnection is damped on the same timescale as the reconnection. (Boozer, 2020) showed that shear Alfvén waves moving along the reconnecting field lines would cause the flattening of the $j_{||}/B$ profile, which leads to the current spike observed in disruptions. The simulations of (Nardon et al., 2023) were consistent with the predicted timescale, and they saw magnetic fluctuations that appeared to be shear Alfvén waves.
@@
(Heyvaerts and Priest, 1983) and (Similon and Sudan, 1989) showed that when shear Alfvén waves propagate along chaotic field lines they are quickly damped. (Huang and Bhattacharjee, 2022) defined reconnection, even in three-coordinate problem, as the transfer of energy from the reconnecting field to the plasma, not the breaking of field line connections. Their definition, which is common in two-coordinate models, led them to the odd conclusion that the chaos of the field lines was irrelevant to reconnection—only the formation of localized currents, which are caused by Alfvén wave damping, are relevant.
@@
The rapid spreading of impurities across the entire plasma is a commonly observed feature of tokamak disruptions. The impurities greatly reduce the damage produced by disruptions by radiating the plasma thermal energy relatively uniformly over the walls rather than having it concentrated in spots where turnstiles formed within the plasma strike the walls as is often the case with runaway electrons. The reason for the impurity spreading did not have a firm theoretical basis, but the Bohm-like transport that is produced when there is an electron pressure gradient along chaotic field lines, (Boozer, 2024a) and Section 4 provide an explanation. The simulations of (Nardon et al., 2023) saw electric fields consistent in magnitude, but the equation in their code which gave the electric field was not simply related to the two-fluid equations that (Boozer, 2024a) used to obtain the Bohm-like transport.
@@
(Breizman et al., 2019) reviewed the problem of the conversion of the plasma current from being carried by near-thermal to being carried by relativistic electrons, which is often viewed as the most dangerous feature of disruptions. Even though the total energy in the relativistic electrons is only about ten percent of the pre-disruption plasma thermal energy, impurity radiation spreads the thermal energy losses almost uniformly over the wall. Relativistic electron losses onto the walls can be extremely concentrated in both space and time, which means they are far more destructive. As pointed out by (Boozer and Punjabi, 2016) the spatial and temporal concentration can be explained by the breaking of the outermost magnetic surface that is confining the runaway electrons in an internal bounded chaotic region. As the outermost surface breaks it forms a cantorus. The slower the turnstiles form compared to the timescale for the relativistic electrons to cover the chaotic region the more tightly collimated are the turnstiles on the cantorus. When the turnstiles form rapidly compared to the timescale for the relativistic electrons to cover the chaotic region, the turnstiles become arbitrary large, which spreads the relativistic electrons so broadly on the walls that they do no damage. The highly damaging case is often observed and brings fear, but (Reux et al., 2021) and (Paz-Soldan et al., 2021) have observed on the JET and the DIII-D tokamaks the benign spreading of the relativistic electrons when the plasma becomes unstable to a rapidly growing magnetic perturbation.
@@
(Boozer, 2025a) has pointed out that the frequency with which disruptions occur in tokamaks may be in large part be due to the far larger poloidal flux produced by the plasma current than the change in the poloidal flux that occurs if the profile of the density of the net plasma current $j_{||}$ is varied over the full range of profiles stable to tearing modes. This flux ratio is approximately a factor of ten. Equation (2), $\partial\psi_{p}/\partial t=V_{\ell}$, implies the plasma will disrupt within a shorter time than the time required to shutdown by removing the plasma-produced poloidal flux or the natural length of a flattop pulse unless a spatially constant loop voltage remains consistent with tearing mode stability. As shown in Equation 19, the topology of the magnetic field is held fixed when $\vec{\nabla}V_{\ell}=0$. An implication is that poloidal flux can be removed or maintained without producing a disruption when
@@
The profile of $j_{||}$ can in principle be controlled by plasma heating and current drive. During the flattop of a fusion pulse, the required power to produce a large change in the $j_{||}$ profile is comparable to the power produced by the alpha-particle heating. This is obvious for direct heating. (Boozer, 1988) showed the power required to drive the full plasma current in a power plant is greater than $j_{cd}E_{ch}$ integrated over the plasma volume where $E_{ch}$ is the Connor-Hastie electric field (Connor and Hastie, 1975). The implication is the power required to drive the total plasma current is comparable to the power produced by alpha-particle heating. Careful control of the injected power is required to properly control the profile of $j_{||}$, but the available diagnostics may not provide adequate information for such control even if it were in principle possible. This disruption issue can be circumvented by the use of stellarator magnetic field to produce the rotational transform by external coil currents rather than by the plasma current.